OpenStudy (anonymous):

What number is 60% of 50?

OpenStudy (anonymous):

"FAN AND MEDAL"

OpenStudy (zarkon):

these questions can be answer by wolfram alpha http://www.wolframalpha.com/input/?i=60%25+of+50

OpenStudy (alexandervonhumboldt2):

muliply the percent by the number

OpenStudy (alexandervonhumboldt2):

60%=60/100 50*60/100=?

OpenStudy (anonymous):

THANKYOU

OpenStudy (alexandervonhumboldt2):

answer is wrong without explanations

OpenStudy (mathmale):

To find a given percentage of a quantity (such as 50), convert the percentage to a decimal fraction and multiply that quantity by it. @ncruz: please convert 60% to a decimal fraction.; Next, mult. 50 by that decimal fraction.

OpenStudy (alexandervonhumboldt2):

Please do not give direct answers. We are not answering servise .

OpenStudy (anonymous):

ok ok go slow

OpenStudy (anonymous):

so first watt do i do

OpenStudy (alexandervonhumboldt2):

convert percent to a faction or decimal

OpenStudy (mathmale):

@alexandervonHumboldt2: Please expand and clarify your response: "answer is wrong without explanations" @ncruz: I explained that in two steps, above. Please look up that post and follow my suggestions.

OpenStudy (alexandervonhumboldt2):

@mathmale i made a statistics. When people give answers without showing work they are 90% wrong.

OpenStudy (anonymous):

0.60 0.50

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

0.60 is correct

OpenStudy (alexandervonhumboldt2):

but where did 0.50 derived from?

OpenStudy (anonymous):

umh

OpenStudy (alexandervonhumboldt2):

we do not convert the number into decimal. anyways you converted it wrong.

OpenStudy (anonymous):

50/100?

OpenStudy (alexandervonhumboldt2):

ok so we have 0.60.' now we multply the number by the decimal: 50*0.60=?

OpenStudy (mathmale):

@AlexandervonHumboldt2 : Respond respectfully to my request, please. This is no place for joking. Several people have been spouting "answers" at the same time, and so it becomes difficult to know which "wrong answer" you're refering to.

OpenStudy (anonymous):

30?

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

i dount understand i got 30 write or rong

OpenStudy (alexandervonhumboldt2):

30 is correct

OpenStudy (anonymous):

ok so thanks avery one how help me

OpenStudy (anonymous):

let me yust fannd you

OpenStudy (alexandervonhumboldt2):

@mathmale i think we are misunderstanding each other. lets stop that conversation